For How Long Must A Current Of 1.5 Ampere Be Passed Through An Aqueous Solution Of A Copper Salt During Electrolysis In Order To Deposit 2.50g Of Copper (cu=63.5)


1
Expert's answer
2022-04-13T15:33:04-0400

Solution:

Current, I = 1.5 A

Mass of copper (Cu), m = 2.50 g

Molar mass of copper (Cu), M = 63.5 g/mol

Cu2+ + 2e → Cu0 (Z = 2)


Faraday's 1st law: The mass of a substance produced at an electrode during electrolysis is proportional to the number of moles of electrons (the quantity of electricity) transferred at that electrode.

Faraday's 1st law can be expressed as:

m = (Q × M) / (Z × F)

where:

m = mass of material (g)

M = molar mass of material (g/mol)

Q = I × t = quantity of electricity (C)

Z = number of electrons

F = Faraday constant, 96485 (C mol−1)


Z × F × m = I × t × M

t = (Z × F × m) / (I × M)

t = (2 × 96485 × 2.50) / (1.5 × 63.5) = 5065 s

t = 5065 s


Answer: 5065 seconds